OpenStudy (anonymous):

can someone help me solve this equation: 3.03(x-7)+x=0.5(3-x)

OpenStudy (anonymous):

\[3.03(x-7)+x=0.5(3-x)\] \[3.03x-21.21+x=1.5-0.5x\] \[4.53x=22.71\] \[x=5.013\]

OpenStudy (anonymous):

x=5.0132

OpenStudy (anonymous):

thank you for your help